OpenStudy (anonymous):

If the equation of the normal to the curve y = 2x2 + 4.5x + 1 is y = −2x + c, find the: a) coordinates at the point of tangency b) value of c.

OpenStudy (anonymous):

take the derivative to get a formula for the slope of the tangent line the normal line has slope as the negative reciprocal of the tangent , so the slope of the tangent line will be \(\frac{1}{2}\) set the derivative equal to \(\frac{1}{2}\) solve for \(x\)

OpenStudy (anonymous):

I've done that now and I tried to continue on from there but the answer I got is different from the book's answer.

OpenStudy (amistre64):

y = 2x2 + 4.5x + 1 y' = 4x + 4.5 since y' = slope of tangent; -1/y' = slope of normal -1/y' = \(\Large -\frac{1}{4x + 4.5}\)

OpenStudy (amistre64):

for what value of x does y' = -2?

OpenStudy (amistre64):

gotta readjust my glasses .. the normal to the curve has the equation: -2x+c

OpenStudy (amistre64):

satellites workings are becoming more clear to me now :)

OpenStudy (amistre64):

for what value of x does y' = 4x + 4.5 = 1/2

OpenStudy (amistre64):

4x + 4.5 = 1/2 8x + 9 = 1 8x = -8

OpenStudy (anonymous):

yes, I also got x = -1 for that step.

OpenStudy (amistre64):

ok, then when x=-1, what does y = for the point we are looking at? y = 2x2 + 4.5x + 1 y = 2 - 4.5 + 1

OpenStudy (anonymous):

oh wait my bad, I missed the plus one at the end because I forgot to write it in the beginning first equation ==, let me try again.

OpenStudy (anonymous):

ok, I got it, thanks so much.

OpenStudy (amistre64):

Ny = -2x +2(-1) + y |---------| = c youre welcome